Grade 12th passMechanics

A ball released from the top of a tower travels 11/36 of the height of the tower in the last second of its journey. the height of the tower is?

Nikita
9 Years agoGrade 12th pass

1 Answer

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
Assuming height of the tower h meter and time t we get,
=>h = 1/2gt^2 -------------------(i) 
Distance travelled in (t-1) secs, 
=>s = 1/2g(t-1)^2 ---------------(ii) 
Subtract both the eqns. we get,
=>11h/36 = 1/2g(2t - 1) ------------(iii) 
 11/36 x 1/2gt^2 = 1/2g(2t-1) 
=>11t^2 = 72t - 36 
=>11t^2 - 72t + 36 = 0 
On sollving further,
=>t = [72+/-60]/22 
=>t = 6 sec OR 0.55 sec 
Therefore. 
h = 176.40 m